# Intuitionistic NF set theory

### v5

@@ 4. Frege cardinals
That completes the proof of the lemma.
5. Order on the cardinals
@@
$\kappa\leq\mu$ ::=:= $\exists a,b\,(a\in\kappa\ \land\ b\in\mu\ \land\ a\subseteq b)$
$\kappa<\mu$ ::=:= $\kappa\leq\mu\ \land\ \kappa\neq\mu.$
@@
$\kappa\leq\mu$ ::=:= $\exists a,b\,(a\in\kappa\ \land\ b\in\mu\ \land\ a\subseteq b\ \land\ b=a\cup(b-a))$
$\kappa<\mu$ ::=:= $\kappa\leq\mu\ \land\ \kappa\neq\mu.$
@@
The image of $a$ under $f$, written $f``a$, is defined by
@@
If $f$ is a function then $f``a$ is the set of values $f(x)$ for $x\in a$.
@@
The image of a separable subset under a similarity is a separable subset. More precisely, let $f:b\to c$ be a similarity and suppose $b=a\cup(b-a)$. Let $e=f``a$ be the image of $a$ under $f$. Then $c=e\cup(c-e)$.
Proof. We have
(31) $e\cup(c-e)$$\subseteq$ $c$
@@
(32) $c$$\subseteq$ $e\cup(c-e)$
since if $q\in c$ then $q=f(p)$ for some $p\in b$, and $p\in a\ \lor\ p\in b-a$, since $b=a\cup(b-a)$, and if $p\in a$ then $q\in e$, while if $p\in b-a$ then $q\in c-e$. Combining (31) and (32), we have $c=e\cup(c-e)$ as desired. That completes the proof of the lemma.
@@
Proof. Straightforward; requires about 75 inferences that we choose to omit here.
@@
Proof. Suppose $\kappa\leq\lambda$ and $\lambda\leq\mu$. We must show $\kappa\leq\mu$. Since $\kappa<\lambda$ and $\lambda<\mu$, there exist $a\in\kappa$, $b,c\in\lambda$, and $d\in\mu$ such that $a\subseteq b$ and $c\subseteq d$, and $b=a\cup(b-a)$, and $d=c\cup(d-c)$. By Lemma 4.9, $b\sim c$, since both belong to $\lambda$. Let $f:b\to c$ be one-to-one and onto. Let $e=f``a$. Then $e\subseteq c$ and $a\sim e$. So $e\in\kappa$, by Lemma 4.8. Then $e\subseteq d$. By Lemma 5.4 we have
(33) $c$ $=$ $e\cup(c-e)$
@@
as desired. Then $\kappa\leq\mu$ as desired. That completes the proof of the lemma.
Lemma 5.7.
(i) If two finite cardinals have a common member, then they are equal.
(ii) Two distinct finite cardinals are disjoint.
Proof. Part (ii) is the contrapositive of (i), so it suffices to prove (i). Let $\kappa$ and $\mu$ belong to ${\mathbb{F}}$. Suppose $x$ belongs to both $\kappa$ and $\mu$. We must show $\kappa=\mu$. By extensionality, it suffices to show that $\kappa$ and $\mu$ have the same members. Let $y\in\kappa$. Then by Lemma 4.9, $y\sim x$. By Lemma 4.8, $y\in\mu$. Therefore $\kappa\subseteq\mu$. Similarly $\mu\subseteq\kappa$. That completes the proof of the lemma.
@@
Proof. Left to right: Suppose $\kappa<\mu$. Then by definition of $<$, $\kappa\leq\mu$ and $\kappa\neq\mu$. By definition of $\leq$, there exist $x$ and $y$ with $x\in\kappa$, $y\in\mu$, and $x\subseteq y$ and $y=x\cup(y-x)$. By Lemma 5.7, which applies because $\kappa\neq\mu$, we have $x\neq y$. Therefore $x\subset y$ as desired. That completes the proof of the left-to-right implication.
Right to left: Suppose $x\in\kappa$ and $y\in\mu$ and $x\subset y$ and $y=x\cup(y-x)$. Then $\kappa\leq\mu$ by definition. We must show $\kappa\neq\mu$. If $\kappa=\mu$ then $y\sim x$, by Lemma 4.9. Then $y$ is similar to a proper subset of $y$, namely $x$. Since $y\in\mu$ and $\mu\in{\mathbb{F}}$, by Lemma 4.4, $y$ is finite. Since $y$ is similar to a proper subset of itself (namely $x$), Theorem 3.24 implies that $y$ is not finite, which is a contradiction. That completes the proof of the lemma.
@@
Proof. Left-to-right. Suppose $\kappa\leq\mu$. Then by definition of $\leq$, there exist $x\in\kappa$ and $y\in\mu$ with $x\subseteq y$ and
(34) $y$ $=$ $x\cup(y-x)$
@@
We have $b\sim y$ by Lemma 4.9. So $y\sim b$. Let $f:y\to b$ be one-to-one and onto. Let $a=f``(x)$. Then $a\subseteq b$ and $x\sim a$. By Lemma 4.8, $a\in\kappa$. By Lemma 5.4 and (34), we have $b=a\cup(b-a)$. That completes the proof of the left-to-right implication.
Right-to-left. Suppose $\forall b\in\mu\,\exists a\in\kappa\,(a\subseteq b\ \land\ b=a\cup(b-a))$. Since $\mu$ is a inhabited, there exists $b\in\mu$. Then $\exists a\in\kappa\,(a\subseteq b\ \land\ b=a\cup(b-a))$. That completes the proof of the lemma.
@@
Proof. Since $x\in\kappa^{+}$, there exists $z\in\kappa$ and $a\not\in z$ such that $x=z\cup\{a\}$. Since $c\in x$, we have $c\in z\ \lor\ c=a$. If $c=a$ then $z=x-\{c\}\in\kappa$ and we are done. Therefore we may assume $c\in z$ and $c\neq a$.
@@
(35) $(z\cup\{a\})-\{c\}=(z-\{c\})\cup\{a\}$
Since $x\in\kappa^{+}$, $x$ is finite, by Lemma 4.4. By Lemma 3.3, $x$ has decidable equality. Then
@@
(36) $(z\cup\{a\})-\{c\}\in\kappa$
Since $z\cup\{a\}=x$, that implies $x-\{c\}\in\kappa$, which is the conclusion of the lemma. That completes the proof of the lemma.
@@
Proof. Left to right. Suppose $\kappa\leq\mu$. Since $\mu^{+}$ is inhabited, there is some $y\in\mu$ and some $c\not\in y$, so $y\cup\{c\}\in\mu^{+}$. By Lemma 5.9, there is a separable subset $x\subseteq y$ with $x\in\kappa$. Then $x\cup\{c\}\in\kappa^{+}$ and $x\cup\{c\}\subset y\cup\{c\}$. We have to show that
(38) $y\cup\{c\}$ $=$ $(x\cup\{c\})\cup(y\cup\{c\}-(x\cup\{c\})).$
@@
Right-to-left. Suppose $\kappa^{+}\leq\mu^{+}$. Then there exist $x\in\kappa^{+}$ and $y\in\mu^{+}$ with $x\subseteq y$ and $y=x\cup(y-x)$. By Lemma 4.15, $x$ is inhabited, so there exists $c\in x$. Since $x\subseteq y$, also $c\in y$. Then by Lemma 5.10, $x-\{c\}\in\kappa$ and $y-\{c\}\in\mu$. Since $y\in\mu^{+}$, $y$ is finite, by Lemma 4.4. By Lemma 3.3, $y$ has decidable equality. Then
(39) $u\in y\ {\rightarrow}\ u=c\ \lor\ u\neq c$
. Since $y=x\cup(y-x)$, we have
(40) $u\in y\ {\rightarrow}\ u\in x\ \lor\ u\not\in x$
@@
Therefore $\kappa\leq\mu$. That completes the proof of the lemma.
@@
For $\lambda$ and $\mu$ in ${\mathbb{F}}$, if $\lambda+$ and $\mu^{+}$ are inhabited, then
@@
Proof. Left to right is immediate. We take up the right to left implication. Suppose $\kappa^{+}=\mu^{+}$. By Lemma 5.7, it suffices to show that $\kappa\cap\mu$ is inhabited. Since $\kappa^{+}$ is inhabited, there exists $y\in\kappa^{+}$. By definition of successor, $y$ has the form $y=x\cup\{a\}$ for some $x\in\kappa$ and $a\not\in x$. We will prove $x\in\mu$. Since $\mu^{+}=\kappa^{+}$ we have $x\cup\{a\}\in\mu^{+}$. Then by Lemma 5.10, $x\cup\{a\}-\{a\}\in\mu$. Since $x\cup\{a\}\in\mu^{+}$, $x\cup\{a\}$ is finite, by Lemma 4.4. By Lemma 3.3, $x\cup\{a\}$ has decidable equality. Then $x\cup\{a\}-\{a\}=x$, so $x\in\mu$. Then $x\in\kappa\cap\mu$ as claimed. That completes the proof of the lemma.
@@
Let $x$ be a separable subset of $y$, that is, $x\subseteq y$ and $y=x\cup(y-x)$. Then $y-x=\Lambda{\ \leftrightarrow\ }y=x$.
Proof. Suppose $x\subseteq y$ and $y=x\cup(y-x)$. Left to right: suppose $y-x=\Lambda$; we must show $y=x$. If $u\in x$ then by $y=x\cup(y-x)$ we have $u\in y$. Conversely, if $u\in y$ then $u\in x\ \lor\ u\not\in x$. If $u\in x$ we are done. If $u\not\in x$ then $u\in y-x$, so $u\in y$. That completes the left-to-right direction. Right to left: Suppose $y=x$. Then $y-x=x-x=\Lambda$. That completes the proof of the lemma.
@@
Proof. Left-to-right. Suppose $\kappa<\mu$. By definition that means $\kappa\leq\mu$ and $\kappa\neq\mu$. By Lemma 5.11, $\kappa^{+}\leq\mu^{+}$. We have to show $\kappa^{+}\neq\mu^{+}$. Suppose $\kappa^{+}=\mu^{+}$. Since $\mu^{+}$ is inhabited, there is an element $y\cup\{c\}$ of $\mu^{+}$ with $y\in\mu$ and $c\not\in y$. Since $\kappa^{+}=\mu^{+}$, we also have $y\cup\{c\}\in\kappa^{+}$. Since $y\in\mu$, by Lemma 4.4, $y$ is finite. Since $\mu^{+}$ is inhabited, $\mu$ is also inhabited. Since $\kappa<\mu$, by Lemma 5.9, there exists a separable subset $x$ of $y$ with $x\in\kappa$. By Lemma 4.4, $x$ is finite. By Lemma 3.20, $y-x$ is finite. Since $\kappa\neq\mu$, we have $x\neq y$, by Lemma 5.7. Then, since $x$ is a separable subset of $y$, $y-x$ is not empty, by Lemma 5.13. Since it is finite, by Lemma 3.4, $y-x$ is inhabited. Hence there exists some $b\in y$ with $b\not\in x$. Then $x\cup\{b\}\in\kappa^{+}$. Then $x\cup\{b\}$ and $y\cup\{c\}$ both belong to $\kappa^{+}$.
@@
Right-to-left: Suppose $\kappa^{+}<\mu^{+}$. Then $\kappa^{+}\leq\mu^{+}$ and $\kappa^{+}\neq\mu^{+}$. By Lemma 5.11, $\kappa\leq\mu$, and since successor is a function, $\kappa\neq\mu$. That completes the proof of the lemma.
@@
Proof. By induction on $\kappa$. More explicitly, define
@@
Then by definition of ${\mathbb{F}}$ (or, if you prefer, “by induction on $\kappa$”), ${\mathbb{F}}\subseteq W$. Then by the definition of union, $\kappa\in{\mathbb{F}}\ {\rightarrow}\ \kappa={\sf zero}\ \lor\ \kappa\neq{\sf zero}$. That completes the proof of the lemma.
@@
Proof. We prove by induction on $\kappa$ that for all $\mu$ we have the assertion in the statement of the lemma. Lemma 4.5 justifies this method of proof. The formula is stratified since the relation $x<y$ is definable.
@@
and exactly one of the three holds. If $\mu\leq{\sf zero}$, then we would have $x\in\mu$ and $x$ a separable subset of $y$ and $y\in{\sf zero}$; but the only member of zero is $\Lambda$, so $x=y=\Lambda$. Then $\Lambda\in\mu$ and $\Lambda\in{\sf zero}$, so by Lemma 5.7, $\mu={\sf zero}$. Thus $\mu<{\sf zero}$ is impossible and $\mu\leq{\sf zero}$ if and only if $\mu={\sf zero}$. If $\mu\in{\mathbb{F}}$ then by Lemma 5.16, $\mu={\sf zero}\lor\mu\neq{\sf zero}$; and if $\mu\neq{\sf zero}$ then ${\sf zero}<\mu$, since $\Lambda$ is a separable subset of any $x\in\mu$.
@@
(42) $\kappa^{+}<\mu\ \lor\ \kappa^{+}=\mu\ \lor\ \mu<\kappa^{+}$
@@
(43) $\kappa^{+}<\lambda^{+}\ \lor\kappa^{+}=\mu^{+}\ \lor\ \mu^{+}<\lambda^{+}.$
@@
and exactly one of the three holds. By Lemma 5.14 and Lemma 5.12, each disjunct is equivalent to one of the disjuncts of (43). That completes the induction step. That completes the proof of the theorem.
@@
Proof. Let $\kappa,\mu\in{\mathbb{F}}$. We must show $\kappa=\mu\ \lor\kappa\neq\mu$. By Theorem 5.17, we have $\kappa<\mu$ or $\kappa=\mu$ or $\mu<\kappa$, and exactly one of the disjuncts holds. Therefore $\kappa\neq\mu$ is equivalent to $\kappa<\mu\ \lor\ \mu<\kappa$. That completes the proof of the corollary.
@@
Proof. Suppose $\kappa\in{\mathbb{F}}$. By Corollary 4.7, $\kappa$ is inhabited. Let $a\in\kappa$. Since $a$ is a separable subset of $a$, we have $\kappa\leq\kappa$ by the definition of $\leq$. That completes the proof.
@@
Proof. Suppose $\kappa,\mu\in{\mathbb{F}}$. By Theorem 5.17 we have $\kappa<\mu\ \lor\ \kappa=\mu\ \lor\ \mu<\kappa$, and exactly one of the three disjuncts holds.
Left to right: Suppose $\kappa\leq\mu$. By Definition 5.2, there exist $a$ and $b$ with $a\in\kappa$, $b\in\mu$, $a\subseteq b$, and $b=a\cup(b-a)$. By Lemma 4.4, $a$ and $b$ are finite. By Lemma 3.20, $b-a$ is finite. By Lemma 3.4, $b-a$ is empty or inhabited.
Case 1, $b-a=\Lambda$. I say $b=a$. By extensionality, it suffices to prove $t\in b{\ \leftrightarrow\ }t\in a$. Left to right: assume $t\in b$. Since $b=a\cup(b-a)$ we have $t\in a\ \lor\ t\in b-a$. But $t\not\in b-a$, since $b-a=\Lambda$. Therefore $t\in a$. Right to left: assume $t\in a$. Since $a\subseteq b$ we have $t\in b$. Therefore $b=a$ as claimed.
@@
Case 2, $b-a$ is inhabited. Then $a$ is a proper subset of $b$. By Lemma 5.8, $\kappa<\mu$. That completes Case 2. That completes the left to right direction.
Right to left: Suppose $\kappa<\mu$. Then by definition of $<$, we have $\kappa\leq\mu$. On the other hand, if $\kappa=\mu$ then $\kappa\leq\mu$ by Lemma 5.19. That completes the proof of the lemma.
@@
Proof. By Lemma 5.20, it suffices to prove
(44) $(\kappa<\mu\ \lor\ \kappa=\mu)\ \land\ (\mu<\kappa\lor\mu=\kappa)\ {\rightarrow}\ \kappa=\mu.$
@@
and exactly one of the three disjuncts holds. Now (44) follows by propositional logic.
@@
Proof. By Lemma 5.6, we have $\kappa\leq\lambda$. We must show $\kappa\neq\lambda$. Suppose $\kappa=\lambda$. Since $\kappa<\mu$ we have $\lambda<\mu$. Hence $\lambda\leq\mu$. By hypothesis $\mu\leq\lambda$. By Lemma 5.21, $\mu=\lambda$, contradicting $\mu<\lambda$. That completes the proof of the lemma.

### v6

@@ 4 Frege cardinals
Lemma 4.24.
(i) If two finite cardinals have a common member, then they are equal.
(ii) Two distinct finite cardinals are disjoint.
Proof.
Part (ii) is the contrapositive of (i), so it suffices to prove (i). Let $\kappa$ and $\mu$ belong to $\mathbb{F}$}. Suppose $x$ belongs to both $\kappa$ and $\mu$. We must show $\kappa=\mu$. By extensionality, it suffices to show that $\kappa$ and $\mu$ have the same members. Let $y\in\kappa$. Then by Lemma 4.9, $y\sim x$. By Lemma 4.8, $y\in\mu$. Therefore $\kappa\subseteq\mu$. Similarly $\mu\subseteq\kappa$. ∎
Lemma 4.25.
Let $x$ and $y$ be finite sets. Then
$x\sim y\ {\rightarrow}\ |x|=|y|.$
Proof.
Assume $x\in{\sf FINITE}$ and $y\in{\sf FINITE}$ and $x\sim y$. Then
$x\in|x|$   by Lemma 4.11
$y\in|y|$   by Lemma 4.11
$|x|\in{\mathbb{F}}$   by Lemma 4.21
$|y|\in{\mathbb{F}}$   by Lemma 4.21
$y\in|x|$   by Lemma 4.8
$|x|=|y|$   by Lemma 4.24
5 Order on the cardinals
@@
$\kappa\leq\mu$ $:=$ $\exists a,b\,(a\in\kappa\ \land\ b\in\mu\ \land\ a\subseteq b)$
$\kappa<\mu$ $:=$ $\kappa\leq\mu\ \land\ \kappa\neq\mu.$
@@
$\kappa\leq\mu$ $:=$ $\exists a,b\,(a\in\kappa\ \land\ b\in\mu\ \land\ a\subseteq b\ \land\ b=a\cup(b-a))$
$\kappa<\mu$ $:=$ $\kappa\leq\mu\ \land\ \kappa\neq\mu.$
@@
The image of $a$ under $f$, written faf``a, is defined by
@@
If $f$ is a function then faf``a is the set of values $f(x)$ for $x\in a$.
@@
The image of a separable subset under a similarity is a separable subset. More precisely, let $f:b\to c$ be a similarity and suppose $b=a\cup(b-a)$. Let e=fae=f``a be the image of $a$ under $f$. Then $c=e\cup(c-e)$.
Proof.
We have
$e\cup(c-e)$ $\subseteq$ $c$ (31)
@@
$c$ $\subseteq$ $e\cup(c-e)$ (32)
since if $q\in c$ then $q=f(p)$ for some $p\in b$, and $p\in a\ \lor\ p\in b-a$, since $b=a\cup(b-a)$, and if $p\in a$ then $q\in e$, while if $p\in b-a$ then $q\in c-e$. Combining (31) and (32), we have $c=e\cup(c-e)$ as desired. ∎
@@
Proof.
Straightforward; requires about 75 inferences that we choose to omit here. ∎
@@
Proof.
Suppose $\kappa\leq\lambda$ and $\lambda\leq\mu$. We must show $\kappa\leq\mu$. Since $\kappa<\lambda$ and $\lambda<\mu$, there exist $a\in\kappa$, $b,c\in\lambda$, and $d\in\mu$ such that $a\subseteq b$ and $c\subseteq d$, and $b=a\cup(b-a)$, and $d=c\cup(d-c)$. By Lemma 4.9, $b\sim c$, since both belong to $\lambda$. Let $f:b\to c$ be one-to-one and onto. Let e=fae=f``a. Then $e\subseteq c$ and $a\sim e$. So $e\in\kappa$, by Lemma 4.8. Then $e\subseteq d$. By Lemma 5.4 we have
$c$ $=$ $e\cup(c-e)$ (33)
@@
as desired. Then $\kappa\leq\mu$ as desired. ∎
@@
Proof.
Left to right: Suppose $\kappa<\mu$. Then by definition of <<, $\kappa\leq\mu$ and $\kappa\neq\mu$. By definition of $\leq$, there exist $x$ and $y$ with $x\in\kappa$, $y\in\mu$, and $x\subseteq y$ and $y=x\cup(y-x)$. By Lemma 4.24, which applies because $\kappa\neq\mu$, we have $x\neq y$. Therefore $x\subset y$ as desired. That completes the proof of the left-to-right implication.
Right to left: Suppose $x\in\kappa$ and $y\in\mu$ and $x\subset y$ and $y=x\cup(y-x)$. Then $\kappa\leq\mu$ by definition. We must show $\kappa\neq\mu$. If $\kappa=\mu$ then $y\sim x$, by Lemma 4.9. Then $y$ is similar to a proper subset of $y$, namely $x$. Since $y\in\mu$ and $\mu\in{\mathbb{F}}$, by Lemma 4.4, $y$ is finite. Since $y$ is similar to a proper subset of itself (namely $x$), Theorem 3.24 implies that $y$ is not finite, which is a contradiction. ∎
@@
Proof.
Left-to-right. Suppose $\kappa\leq\mu$. Then by definition of $\leq$, there exist $x\in\kappa$ and $y\in\mu$ with $x\subseteq y$ and
$y$ $=$ $x\cup(y-x)$ (34)
@@
We have $b\sim y$ by Lemma 4.9. So $y\sim b$. Let $f:y\to b$ be one-to-one and onto. Let a=f(x)a=f``(x). Then $a\subseteq b$ and $x\sim a$. By Lemma 4.8, $a\in\kappa$. By Lemma 5.4 and (34), we have $b=a\cup(b-a)$. That completes the proof of the left-to-right implication.
Right-to-left. Suppose $\forall b\in\mu\,\exists a\in\kappa\,(a\subseteq b\ \land\ b=a\cup(b-a))$. Since $\mu$ is a inhabited, there exists $b\in\mu$. Then $\exists a\in\kappa\,(a\subseteq b\ \land\ b=a\cup(b-a))$. ∎
@@
Proof.
Since $x\in\kappa^{+}$, there exists $z\in\kappa$ and $a\not\in z$ such that $x=z\cup\{a\}$. Since $c\in x$, we have $c\in z\ \lor\ c=a$. If $c=a$ then $z=x-\{c\}\in\kappa$ and we are done. Therefore we may assume $c\in z$ and $c\neq a$.
@@
$(z\cup\{a\})-\{c\}=(z-\{c\})\cup\{a\}$ (35)
Since $\in\kappa^{+}, xx$ is finite, by Lemma 4.4. By Lemma 3.3, $x$ has decidable equality. Then
@@
$(z\cup\{a\})-\{c\}\in\kappa$ (36)
Since $z\cup\{a\}=x$, that implies $x-\{c\}\in\kappa$, which is the conclusion of the lemma. ∎
@@
Proof.
Left to right. Suppose $\kappa\leq\mu$. Since $\mu^{+}$ is inhabited, there is some $y\in\mu$ and some $c\not\in y$, so $y\cup\{c\}\in\mu^{+}$. By Lemma 5.8, there is a separable subset $x\subseteq y$ with $x\in\kappa$. Then $x\cup\{c\}\in\kappa^{+}$ and $x\cup\{c\}\subset y\cup\{c\}$. We have to show that
$y\cup\{c\}$ $=$ $(x\cup\{c\})\cup(y\cup\{c\}-(x\cup\{c\})).$ (37)
@@
Right-to-left. Suppose $\kappa^{+}\leq\mu^{+}$. Then there exist $x\in\kappa^{+}$ and $y\in\mu^{+}$ with $x\subseteq y$ and $y=x\cup(y-x)$. By Lemma 4.15, $x$ is inhabited, so there exists $c\in x$. Since $x\subseteq y$, also $c\in y$. Then by Lemma 5.9, $x-\{c\}\in\kappa$ and $y-\{c\}\in\mu$. Since $\in\mu^{+}, yy$ is finite, by Lemma 4.4. By Lemma 3.3, $y$ has decidable equality. Then
$u\in y\ {\rightarrow}\ u=c\ \lor\ u\neq c$ (38)
Since $y=x\cup(y-x)$, we have
$u\in y\ {\rightarrow}\ u\in x\ \lor\ u\not\in x$ (39)
@@
Therefore $\kappa\leq\mu$. ∎
@@
For $\lambda$ and $\mu$ in $\mathbb{F}$}, if $\lambda^{+}$ and $\mu^{+}$ are inhabited, then
@@
Proof.
Left to right is immediate. We take up the right to left implication. Suppose $\kappa^{+}=\mu^{+}$. By Lemma 4.24, it suffices to show that $\kappa\cap\mu$ is inhabited. Since $\kappa^{+}$ is inhabited, there exists $y\in\kappa^{+}$. By definition of successor, $y$ has the form $y=x\cup\{a\}$ for some $x\in\kappa$ and $a\not\in x$. We will prove $x\in\mu$. Since $\mu^{+}=\kappa^{+}$ we have $x\cup\{a\}\in\mu^{+}$. Then by Lemma 5.9, $x\cup\{a\}-\{a\}\in\mu$. Since $x\cup\{a\}\in\mu^{+}$, $x\cup\{a\}$ is finite, by Lemma 4.4. By Lemma 3.3, $x\cup\{a\}$ has decidable equality. Then $x\cup\{a\}-\{a\}=x$, so $x\in\mu$. Then $x\in\kappa\cap\mu$ as claimed. ∎
@@
Let $x$ be a separable subset of $y$, that is, $x\subseteq y$ and $y=x\cup(y-x)$. Then $y-x=\emptyset{\ \leftrightarrow\ }y=x$.
Proof.
Suppose $x\subseteq y$ and $y=x\cup(y-x)$. Left to right: suppose $y-x=\emptyset$; we must show $y=x$. If $u\in x$ then by $y=x\cup(y-x)$ we have $u\in y$. Conversely, if $u\in y$ then $u\in x\ \lor\ u\not\in x$. If $u\in x$ we are done. If $u\not\in x$ then $u\in y-x$, so $u\in y$. That completes the left-to-right direction. Right to left: Suppose $y=x$. Then $y-x=x-x=\emptyset$. ∎
@@
Proof.
Left-to-right. Suppose $\kappa<\mu$. By definition that means $\kappa\leq\mu$ and $\kappa\neq\mu$. By Lemma 5.10, $\kappa^{+}\leq\mu^{+}$. We have to show $\kappa^{+}\neq\mu^{+}$. Suppose $\kappa^{+}=\mu^{+}$. Since $\mu^{+}$ is inhabited, there is an element $y\cup\{c\}$ of $\mu^{+}$ with $y\in\mu$ and $c\not\in y$. Since $\kappa^{+}=\mu^{+}$, we also have $y\cup\{c\}\in\kappa^{+}$. Since $y\in\mu$, by Lemma 4.4, $y$ is finite. Since $\mu^{+}$ is inhabited, $\mu$ is also inhabited. Since $\kappa<\mu$, by Lemma 5.8, there exists a separable subset $x$ of $y$ with $x\in\kappa$. By Lemma 4.4, $x$ is finite. By Lemma 3.20, yxy-x is finite. Since $\kappa\neq\mu$, we have $x\neq y$, by Lemma 4.24. Then, since $x$ is a separable subset of $y$, yxy-x is not empty, by Lemma 5.12. Since it is finite, by Lemma 3.4, yxy-x is inhabited. Hence there exists some $b\in y$ with $b\not\in x$. Then $x\cup\{b\}\in\kappa^{+}$. Then $x\cup\{b\}$ and $y\cup\{c\}$ both belong to $\kappa^{+}$.
@@
Right-to-left: Suppose $\kappa^{+}<\mu^{+}$. Then $\kappa^{+}\leq\mu^{+}$ and $\kappa^{+}\neq\mu^{+}$. By Lemma 5.10, $\kappa\leq\mu$, and since successor is a function, $\kappa\neq\mu$. ∎
@@
Proof.
By induction on $\kappa$. More explicitly, define
@@
Then by definition of $\mathbb{F}$} (or, if you prefer, “by induction on $\kappa$”), ${\mathbb{F}}\subseteq W$. Then by the definition of union, $\kappa\in{\mathbb{F}}\ {\rightarrow}\ \kappa={\sf zero}\ \lor\ \kappa\neq{\sf zero}$. ∎
@@
Proof.
We prove by induction on $\kappa$ that for all $\mu$ we have the assertion in the statement of the lemma. Lemma 4.5 justifies this method of proof. The formula is stratified since the relation $x<y$ is definable.
@@
and exactly one of the three holds. If $\mu\leq{\sf zero}$, then we would have $x\in\mu$ and $x$ a separable subset of $y$ and $y\in{\sf zero}$; but the only member of zero is $\emptyset$, so $x=y=\emptyset$. Then $\emptyset\in\mu$ and $\emptyset\in{\sf zero}$, so by Lemma 4.24, $\mu={\sf zero}$. Thus $\mu<{\sf zero}$ is impossible and $\mu\leq{\sf zero}$ if and only if $\mu={\sf zero}$. If $\mu\in{\mathbb{F}}$ then by Lemma 5.15, $\mu={\sf zero}\lor\mu\neq{\sf zero}$; and if $\mu\neq{\sf zero}$ then ${\sf zero}<\mu$, since $\emptyset$ is a separable subset of any $x\in\mu$.
@@
$\kappa^{+}<\mu\ \lor\ \kappa^{+}=\mu\ \lor\ \mu<\kappa^{+}$ (41)
@@
$\kappa^{+}<\lambda^{+}\ \lor\kappa^{+}=\mu^{+}\ \lor\ \mu^{+}<\lambda^{+}.$ (42)
@@
and exactly one of the three holds. By Lemma 5.13 and Lemma 5.11, each disjunct is equivalent to one of the disjuncts of (42). That completes the induction step. ∎
@@
Proof.
Let $\kappa,\mu\in{\mathbb{F}}$. We must show $\kappa=\mu\ \lor\kappa\neq\mu$. By Theorem 5.16, we have $\kappa<\mu$ or $\kappa=\mu$ or $\mu<\kappa$, and exactly one of the disjuncts holds. Therefore $\kappa\neq\mu$ is equivalent to $\kappa<\mu\ \lor\ \mu<\kappa$. ∎
@@
Proof.
Suppose $\kappa\in{\mathbb{F}}$. By Corollary 4.7, $\kappa$ is inhabited. Let $a\in\kappa$. Since $a$ is a separable subset of $a$, we have $\kappa\leq\kappa$ by the definition of $\leq$. ∎
@@
Proof.
Suppose $\kappa,\mu\in{\mathbb{F}}$. By Theorem 5.16 we have $\kappa<\mu\ \lor\ \kappa=\mu\ \lor\ \mu<\kappa$, and exactly one of the three disjuncts holds.
Left to right: Suppose $\kappa\leq\mu$. By Definition 5.2, there exist $a$ and $b$ with $a\in\kappa$, $b\in\mu$, $a\subseteq b$, and $b=a\cup(b-a)$. By Lemma 4.4, $a$ and $b$ are finite. By Lemma 3.20, bab-a is finite. By Lemma 3.4, bab-a is empty or inhabited.
Case 1, $b-a=\emptyset$. I say $b=a$. By extensionality, it suffices to prove $t\in b{\ \leftrightarrow\ }t\in a$. Left to right: assume $t\in b$. Since $b=a\cup(b-a)$ we have $t\in a\ \lor\ t\in b-a$. But $t\not\in b-a$, since $b-a=\emptyset$. Therefore $t\in a$. Right to left: assume $t\in a$. Since $a\subseteq b$ we have $t\in b$. Therefore $b=a$ as claimed.
@@
Case 2, bab-a is inhabited. Then $a$ is a proper subset of $b$. By Lemma 5.7, $\kappa<\mu$. That completes Case 2. That completes the left to right direction.
Right to left: Suppose $\kappa<\mu$. Then by definition of <<, we have $\kappa\leq\mu$. On the other hand, if $\kappa=\mu$ then $\kappa\leq\mu$ by Lemma 5.18. ∎
@@
Proof.
By Lemma 5.19, it suffices to prove
$(\kappa<\mu\ \lor\ \kappa=\mu)\ \land\ (\mu<\kappa\lor\mu=\kappa)\ {\rightarrow}\ \kappa=\mu.$ (43)
@@
and exactly one of the three disjuncts holds. Now (43) follows by propositional logic. ∎
We next prove two variations on trichotomy that are frequently useful.
@@
Proof.
By Lemma 5.6, we have $\kappa\leq\lambda$. We must show $\kappa\neq\lambda$. Suppose $\kappa=\lambda$. Since $\kappa<\mu$ we have $\lambda<\mu$. Hence $\lambda\leq\mu$. By hypothesis $\mu\leq\lambda$. By Lemma 5.20, $\mu=\lambda$, contradicting $\mu<\lambda$. ∎
Lemma 5.22.
Let $\kappa,\mu\in{\mathbb{F}}$. Then
$\kappa<\mu\lor\mu\leq\kappa.$
Proof.
$\kappa<\mu\ \lor\ \kappa=\mu\ \lor\ \mu<\kappa$   by Theorem 5.16 (44)
Case 1, $\kappa<\mu$. Then we are done.
Case 2, $\kappa=\mu$. Then $\kappa\leq\mu$ by Lemma 5.18.
Case 3, $\mu<\kappa$. Then $\mu\leq\kappa$ by the definition of <<. ∎
Lemma 5.23.
Let $\kappa,\mu\in{\mathbb{F}}$. Then
$\kappa\leq\mu\lor\mu<\kappa.$
Proof.
$\kappa<\mu\ \lor\ \kappa=\mu\ \lor\ \mu<\kappa$   by Theorem 5.16 (45)
Case 1, $\kappa<\mu$. Then $\kappa\leq\mu$ by the definition of <<.
Case 2, $\kappa=\mu$. Then $\kappa\leq\mu$ by Lemma 5.18.
Case 3, $\mu<\kappa$. Then we are done. ∎